# Learning Protein-Ligand Binding in Hyperbolic Space

### v1

@@ Abstract
Protein-ligand binding prediction is central to virtual screening and affinity ranking, two fundamental tasks in drug discovery. While recent retrieval-based methods embed ligands and protein pockets into Euclidean space for similarity-based search, the geometry of Euclidean embeddings often fails to capture the hierarchical structure and fine-grained affinity variations intrinsic to molecular interactions. In this work, we propose HypSeek, a hyperbolic representation learning framework that embeds ligands, protein pockets, and sequences into Lorentz-model hyperbolic space. By leveraging the exponential geometry and negative curvature of hyperbolic space, HypSeek enables expressive, affinity-sensitive embeddings that can effectively model both global activity and subtle functional differences-particularly in challenging cases such as activity cliffs, where structurally similar ligands exhibit large affinity gaps. Our mode unifies virtual screening and affinity ranking in a single framework, introducing a protein-guided three-tower architecture to enhance representational structure. HypSeek improves early enrichment in virtual screening on DUD-E from 42.63 to 51.44 (+20.7%) and affinity ranking correlation on JACS from 0.5774 to 0.7239 (+25.4%), demonstrating the benefits of hyperbolic geometry across both tasks and highlighting its potential as a powerful inductive bias for protein-ligand modeling.
@@ 1 Introduction
Modeling protein–ligand interactions is critical for drug discovery, where accurate binding affinity prediction underpins both large-scale virtual screening and fine-grained ligand prioritization. Virtual screening seeks to identify molecules likely to bind a given protein target from large compound libraries, often containing millions or even billions of candidates. Approaches such as molecular docking Friesner et al. (2004); Trott & Olson (2010) estimate binding compatibility by sampling ligand poses and scoring them with physics-based functions. While effective in small-scale settings, these methods are computationally intensive and scale poorly to modern library sizes. Unlike virtual screening, which emphasizes identifying likely binders from vast libraries, affinity ranking focuses on ordering a smaller set of candidate ligands by predicted binding strength, with physics-based techniques like free energy perturbation (FEP+) Wang et al. (2015) offering high accuracy at the cost of extensive molecular dynamics simulations. These limitations restrict the practicality of traditional methods in early-stage drug discovery pipelines.
A notable shift in virtual screening came with DrugCLIP Gao et al. (2023a), which reframed the task as a dense retrieval problem. Rather than predicting binding affinity or docking poses, DrugCLIP learns contrastive embeddings of ligands and protein pockets such that interacting pairs are close in a shared Euclidean space. This design enables efficient similarity-based retrieval and allows for scalable screening across billion-scale compound libraries. Despite its promising performance and efficiency, DrugCLIP struggles to capture fine-grained interaction patterns which are essential for downstream affinity ranking. Recently, LigUnity Feng et al. (2025) extends the retrieval-based framework by unifying virtual screening and affinity ranking into a single training objective. It combines contrastive learning for global interaction patterns with listwise ranking to model pocket-specific ligand preferences, aiming to jointly learn both binding likelihood and relative affinity within a unified embedding space.
@@
We evaluate HypSeek across both large-scale virtual screening and fine-grained affinity ranking tasks. On the DUD-E Mysinger et al. (2012) benchmark, HypSeek improves EF1% from 42.63 to 51.44 (+20.7%), demonstrating strong retrieval performance across targets. For affinity ranking, it increases Spearman correlation on the JACS Wang et al. (2015) dataset from 0.5774 to 0.7239 (+25.4%), consistently outperforming Euclidean baselines. These results highlight the benefits of hyperbolic geometry in capturing both global activity and nuanced affinity variation within a unified embedding space.
@@ 2 Related Work
Virtual Screening. Structure‐based virtual screening traditionally relies on molecular docking methods such as Glide Friesner et al. (2004) and AutoDock Trott & Olson (2010), which predict ligand binding poses and evaluate affinities using physics‐based scoring functions. Some predict binding affinity directly from protein–ligand complex structures by learning scoring functions McNutt et al. (2021b); Jiang et al. (2021); Shen et al. (2022); Cao et al. (2024), while others infer interactions from raw structural inputs Lu et al. (2022); Zhang et al. (2023). A major shift occurred with DrugCLIP Gao et al. (2023a), which introduced contrastive retrieval by aligning ligand and pocket embeddings in a shared Euclidean space for billion‐scale similarity search. This paradigm has since inspired a range of efficient retrieval methods. For example, DrugHash Han et al. (2025a) employs binary hash codes for efficient retrieval with reduced memory cost, and LigUnity Feng et al. (2025) integrates listwise ranking with contrastive screening.
@@
Affinity Ranking. Accurate ranking of ligand binding affinities is essential for lead optimization but remains computationally challenging. Physics‐based methods such as FEP+ (Wang et al., 2015) and MM‐GB/SA (Genheden & Ryde, 2015) deliver high accuracy via alchemical free‐energy calculations and implicit solvent models, respectively, yet they require extensive molecular dynamics sampling. Recent deep learning approaches seek to reduce this cost: PBCNet (Yu et al., 2023) models pairwise ligand differences with graph neural networks, EHIGN (Yang et al., 2024) encodes heterogeneous protein–ligand interaction graphs, and LigUnity (Feng et al., 2025) combines contrastive screening with listwise ranking to jointly address global retrieval and local prioritization.
@@
Hyperbolic Representation Learning. Hyperbolic space has emerged as a powerful embedding manifold for data with latent hierarchical or tree-like structure, owing to its exponential volume growth that preserves hierarchy with low distortion Nickel & Kiela (2017); Chamberlain et al. (2017). Early works demonstrated that embedding taxonomies or graphs in Poincaré or Lorentz models captures hierarchical relations more faithfully than Euclidean counterparts Ganea et al. (2018); Bécigneul & Ganea (2018). This theoretical appeal led to specialized optimization methods and the design of hyperbolic neural layers, including Riemannian gradient algorithms Bonnabel (2013); Bécigneul & Ganea (2019) and Hyperbolic Neural Networks Shimizu et al. (2021), as well as adaptations of convolutional, attention, and graph architectures Gulcehre et al. (2019); Bdeir et al. (2024). Hyperbolic embeddings have demonstrated strong performance across diverse modalities—knowledge graphs and recommender systems Liu et al. (2019); Wang et al. (2021), vision tasks Mettes et al. (2024) such as classification and few-shot learning Khrulkov et al. (2020b); Franco et al. (2023); Liu et al. (2025), and language modeling Dhingra et al. (2018); Tifrea et al. (2019). Recent studies further explore multimodal training in hyperbolic space for vision–language models to capture hierarchical semantics Desai et al. (2023b); Pal et al. (2024); Poppi et al. (2025). Our work is the first to bring hyperbolic space to protein–ligand retrieval, leveraging its inductive bias to separate fine‐grained affinity differences.
@@ 3 Preliminaries
We perform all representation learning in an $n$-dimensional hyperbolic space of constant negative curvature, using the Lorentz model Nickel & Kiela (2018); Lin et al. (2023a); Desai et al. (2023b). This choice affords numerical stability and readily supports geodesic and exponential‐map operations.
@@
where $\|\mathbf{v}\|_{\mathbb{L}}=\sqrt{\langle\mathbf{v},\mathbf{v}\rangle_{\mathbb{L}}}$. In practice, we interpret the output of a Euclidean encoder as a vector in the tangent space at the point $\mathbf{0}=(\frac{1}{\sqrt{\kappa}},0,\dots,0)^{\top}$ on the hyperboloid, and then apply the exponential map $\exp_{\mathbf{0}}^{\kappa}$ to lift it onto $\mathbb{L}^{n}$ Khrulkov et al. (2020a).
@@ 4.2 Multimodal Encoding and Lorentz Mapping
Let $x^{p}$ and $x^{m}$ denote the atom-based inputs (coordinates and types) for a protein pocket and ligand, respectively, and let $S=(s_{1},\dots,s_{L})$ denote the amino acid sequence of a target protein. We define three encoder functions: $g_{\phi}$ and $f_{\theta}$ as SE(3)-equivariant 3D graph transformers for pockets and ligands (following DrugCLIP (Gao et al., 2023a)), and $h_{\psi}$ as a protein sequence encoder based on ESM-2 (Lin et al., 2023b). As illustrated in Figure 2, each encoder maps its input to a vector in $\mathbb{R}^{d_{\mathrm{euc}}}$:
@@
We retain the in‐batch contrastive retrieval losses of DrugCLIP (Gao et al., 2023a) and LigUnity’s listwise ranking term (Feng et al., 2025), applied to the spatial components $\tilde{\mathbf{h}}_{u}=\mathbf{h}_{u}[1:]$. For each assay $A_{i}$ with query modality $u\in\{\mathrm{poc},\mathrm{seq}\}$ and its $B$ candidate ligands $\{v_{j}\}$, we compute similarity logits $s_{i,j}\;=\;\frac{1}{\tau}\,\bigl{\langle}\tilde{\mathbf{h}}_{u_{i}},\tilde{\mathbf{h}}_{v_{j}}\bigr{\rangle}.$
@@
$\mathcal{L}_{\mathrm{l}\to\mathrm{p}}^{(i)}$ $=-\frac{1}{|L_{i}|}\sum_{k\in L_{i}}\log\frac{\exp(s_{k,i})}{\sum_{n=1}^{B}\exp(s_{n,i})},$ (9)
@@
For each assay $A_{i}$ the screened ligands are sorted by measured affinity, yielding an ordered list $(v_{i,1},\dots,v_{i,B})$. Following the Plackett–Luce model Cao et al. (2007), the probability of selecting ligand $v_{i,k}$ at step $k$ (from the remaining set $\mathcal{R}_{i,k}=\{k,k+1,\dots,B\}$) is
@@ 4.4 Hyperbolic Geometry as a Structural Prior
Each pocket defines a surface of admissible directions. Its half-aperture angle is formulated by Le et al. (2019); Desai et al. (2023a) as
@@
Bucket 0 therefore collects the strongest binders and bucket $K$ the weakest. For every ligand we derive a bucket–specific radial limit $r_{i,j}$ and angular–scaling factor $\eta_{i,j}$
@@
where $r_{0}$ and $\eta_{0}$ are the base radius/angle for the strongest tier, and $\Delta r,\Delta\eta>0$ are the per‑tier increments. Smaller $b_{i,j}$ thus yields both a smaller radius cap and a narrower cone. We penalise violations in radius and angle:
@@ 4.5 Addressing Activity Cliffs with Hyperbolic Geometry
While structurally similar ligands often cluster in Euclidean space, such geometry can underrepresent functional differences—especially in activity cliffs, where minor structural changes lead to large affinity shifts. As formalized in Proposition 1, hyperbolic space provides exponentially greater separation via angular variation, offering a principled mechanism for distinguishing such cases. The theoretical derivation is provided in Appendix A.
@@ 4.6 Training and Inference
$\mathcal{L}_{\text{total}}=\underbrace{\alpha_{\mathrm{poc}}\Bigl{(}\mathcal{L}_{\text{cont}}^{\mathrm{poc}\!\leftrightarrow\!\mathrm{lig}}+\lambda_{\text{rank}}\,\mathcal{L}_{\text{rank}}^{\mathrm{poc}}\Bigr{)}}_{\text{pocket}\;\leftrightarrow\;\text{ligand}}+\underbrace{\alpha_{\mathrm{seq}}\Bigl{(}\mathcal{L}_{\text{cont}}^{\mathrm{seq}\!\leftrightarrow\!\mathrm{lig}}+\lambda_{\text{rank}}\,\mathcal{L}_{\text{rank}}^{\mathrm{seq}}\Bigr{)}}_{\text{sequence}\;\leftrightarrow\;\text{ligand}}+\underbrace{\gamma_{\text{cone}}\,\mathcal{L}_{\text{cone}}+\lambda_{\text{ang}}\,R_{\text{ang}}+\lambda_{\text{het}}\,R_{\text{het}}}_{\text{pocket}\;\leftrightarrow\;\text{ligand}}.$ (22)
@@ 5.1 Experimental Settings
Implemention Details. We adopt the same curated assay–level training dataset as LigUnity Feng et al. (2025), which is constructed from ChEMBL (Mendez et al., 2018), BindingDB Gilson et al. (2015), and PDBBind (Liu et al., 2017). For virtual screening, we strictly exclude any target UniProt IDs present in the DUD-E Mysinger et al. (2012), LIT-PCBA Tran-Nguyen et al. (2020) test sets. For affinity ranking tasks, we perform ligand‐level deduplication by removing redundant small molecules and non‐redundant assay IDs. Training is run on four NVIDIA A100 GPUs for 50 epochs,, using the Adam optimizer with an initial learning rate of $1\times 10^{-4}$ and the curvature parameter $\kappa$ (absolute value of negative curvature) fixed to $1$.
Benchmark. In virtual screening, evaluations are performed on DUD-E Mysinger et al. (2012) and LIT-PCBA Tran-Nguyen et al. (2020). DUD-E includes 102 protein targets, each associated with experimentally verified actives and 50 property-matched decoys, designed to test enrichment capability under artificially constructed decoy scenarios. LIT-PCBA, in contrast, contains 15 targets with over 400K experimentally confirmed inactives, offering a more realistic and challenging setting without synthetic decoy bias. For affinity ranking, the evaluation is conducted on JACS Wang et al. (2015) and Merck Schindler et al. (2020). JACS consists of eight high-quality congeneric series extracted from real lead optimization projects, emphasizing precise ranking within narrow chemical series, while Merck serves as a large-scale benchmark for FEP-based lead optimization with diverse chemical scaffolds and higher experimental noise.

### v2

@@ Abstract
Protein-ligand binding prediction is central to virtual screening and affinity ranking, two fundamental tasks in drug discovery. While recent retrieval-based methods embed ligands and protein pockets into Euclidean space for similarity-based search, the geometry of Euclidean embeddings often fails to capture the hierarchical structure and fine-grained affinity variations intrinsic to molecular interactions. In this work, we propose HypSeek, a hyperbolic representation learning framework that embeds ligands, protein pockets, and sequences into Lorentz-model hyperbolic space. By leveraging the exponential geometry and negative curvature of hyperbolic space, HypSeek enables expressive, affinity-sensitive embeddings that can effectively model both global activity and subtle functional differences-particularly in challenging cases such as activity cliffs, where structurally similar ligands exhibit large affinity gaps. Our mode unifies virtual screening and affinity ranking in a single framework, introducing a protein-guided three-tower architecture to enhance representational structure. HypSeek improves early enrichment in virtual screening on DUD-E from 42.63 to 51.44 (+20.7%) and affinity ranking correlation on JACS from 0.5774 to 0.7239 (+25.4%), demonstrating the benefits of hyperbolic geometry across both tasks and highlighting its potential as a powerful inductive bias for protein-ligand modeling. Our code is publicly available at https://github.com/jianhuiwemi/HypSeek.
@@ 1 Introduction
Modeling protein–ligand interactions is critical for drug discovery, where accurate binding affinity prediction underpins both large-scale virtual screening and fine-grained ligand prioritization. Virtual screening seeks to identify molecules likely to bind a given protein target from large compound libraries, often containing millions or even billions of candidates. Approaches such as molecular docking friesner_glide_2004; trott_autodock_vina_2010 estimate binding compatibility by sampling ligand poses and scoring them with physics-based functions. While effective in small-scale settings, these methods are computationally intensive and scale poorly to modern library sizes. Unlike virtual screening, which emphasizes identifying likely binders from vast libraries, affinity ranking focuses on ordering a smaller set of candidate ligands by predicted binding strength, with physics-based techniques like free energy perturbation (FEP+) wang2015accurate offering high accuracy at the cost of extensive molecular dynamics simulations. These limitations restrict the practicality of traditional methods in early-stage drug discovery pipelines.
A notable shift in virtual screening came with DrugCLIP gao2023drugclip, which reframed the task as a dense retrieval problem. Rather than predicting binding affinity or docking poses, DrugCLIP learns contrastive embeddings of ligands and protein pockets such that interacting pairs are close in a shared Euclidean space. This design enables efficient similarity-based retrieval and allows for scalable screening across billion-scale compound libraries. Despite its promising performance and efficiency, DrugCLIP struggles to capture fine-grained interaction patterns which are essential for downstream affinity ranking. Recently, LigUnity feng2025foundation extends the retrieval-based framework by unifying virtual screening and affinity ranking into a single training objective. It combines contrastive learning for global interaction patterns with listwise ranking to model pocket-specific ligand preferences, aiming to jointly learn both binding likelihood and relative affinity within a unified embedding space.
@@
We evaluate HypSeek across both large-scale virtual screening and fine-grained affinity ranking tasks. On the DUD-E mysinger2012dude benchmark, HypSeek improves EF1% from 42.63 to 51.44 (+20.7%), demonstrating strong retrieval performance across targets. For affinity ranking, it increases Spearman correlation on the JACS wang2015accurate dataset from 0.5774 to 0.7239 (+25.4%), consistently outperforming Euclidean baselines. These results highlight the benefits of hyperbolic geometry in capturing both global activity and nuanced affinity variation within a unified embedding space.
@@ 2 Related Work
Virtual Screening. Structure‐based virtual screening traditionally relies on molecular docking methods such as Glide friesner_glide_2004 and AutoDock trott_autodock_vina_2010, which predict ligand binding poses and evaluate affinities using physics‐based scoring functions. Some predict binding affinity directly from protein–ligand complex structures by learning scoring functions mcnutt_gnina_2021; shen_rtmscore_2022; cao_equiscore_2024, while others infer interactions from raw structural inputs lu_tankbind_2022; zhang_karmadock_2023. A major shift occurred with DrugCLIP gao2023drugclip, which introduced contrastive retrieval by aligning ligand and pocket embeddings in a shared Euclidean space for billion‐scale similarity search. This paradigm has since inspired a range of efficient retrieval methods. For example, DrugHash HanHongLi2025 employs binary hash codes for efficient retrieval with reduced memory cost, and LigUnity feng2025foundation integrates listwise ranking with contrastive screening.
@@
Affinity Ranking. Accurate ranking of ligand binding affinities is essential for lead optimization but remains computationally challenging. Physics‐based methods such as FEP+ (wang2015accurate) and MM‐GB/SA (genheden2015mm) deliver high accuracy via alchemical free‐energy calculations and implicit solvent models, respectively, yet they require extensive molecular dynamics sampling. Recent deep learning approaches seek to reduce this cost: PBCNet (yu2023computing) models pairwise ligand differences with graph neural networks, EHIGN (yang2024interaction) encodes heterogeneous protein–ligand interaction graphs, and LigUnity (feng2025foundation) combines contrastive screening with listwise ranking to jointly address global retrieval and local prioritization.
@@
Hyperbolic Representation Learning. Hyperbolic space has emerged as a powerful embedding manifold for data with latent hierarchical or tree-like structure, owing to its exponential volume growth that preserves hierarchy with low distortion nickel2017poincare; chamberlain2017neural. Early works demonstrated that embedding taxonomies or graphs in Poincaré or Lorentz models captures hierarchical relations more faithfully than Euclidean counterparts ganea2018hyperbolic; becigneul2018riemannian. This theoretical appeal led to specialized optimization methods and the design of hyperbolic neural layers, including Riemannian gradient algorithms Bonnabel13; BecigneulG19 and Hyperbolic Neural Networks shimizu2021hyperbolicneuralnetworks, as well as adaptations of convolutional, attention, and graph architectures gulcehre2018hyperbolicattentionnetworks; bdeir2024fully. Hyperbolic embeddings have demonstrated strong performance across diverse modalities—knowledge graphs and recommender systems liu2019hyperbolicgraphneuralnetworks; wang2021fully, vision tasks MettesGhadimiAtighKellerRessel2024 such as classification and few-shot learning khrulkov2020hyperbolic; franco2023hyperbolic; liu2025hyperboliccategorydiscovery, and language modeling dhingra2018embedding; TifreaBG19. Recent studies further explore multimodal training in hyperbolic space for vision–language models to capture hierarchical semantics desai2023hyperbolic; pal2024compositional; poppi2025hyperbolic. Our work is the first to bring hyperbolic space to protein–ligand retrieval, leveraging its inductive bias to separate fine‐grained affinity differences.
@@ 3 Preliminaries
We perform all representation learning in an $n$-dimensional hyperbolic space of constant negative curvature, using the Lorentz model nickel2018learning; lin2023hyperbolic; desai2023hyperbolic. This choice affords numerical stability and readily supports geodesic and exponential‐map operations.
@@
where $\|\mathbf{v}\|_{\mathbb{L}}=\sqrt{\langle\mathbf{v},\mathbf{v}\rangle_{\mathbb{L}}}$. In practice, we interpret the output of a Euclidean encoder as a vector in the tangent space at the point $\mathbf{0}=(\frac{1}{\sqrt{\kappa}},0,\dots,0)^{\top}$ on the hyperboloid, and then apply the exponential map $\exp_{\mathbf{0}}^{\kappa}$ to lift it onto $\mathbb{L}^{n}$ Khrulkov2020.
@@ 4.2 Multimodal Encoding and Lorentz Mapping
Let $x^{p}$ and $x^{m}$ denote the atom-based inputs (coordinates and types) for a protein pocket and ligand, respectively, and let $S=(s_{1},\dots,s_{L})$ denote the amino acid sequence of a target protein. We define three encoder functions: $g_{\phi}$ and $f_{\theta}$ as SE(3)-equivariant 3D graph transformers for pockets and ligands (following DrugCLIP (gao2023drugclip)), and $h_{\psi}$ as a protein sequence encoder based on ESM-2 (lin2023evolutionary). As illustrated in Figure 2, each encoder maps its input to a vector in $\mathbb{R}^{d_{\mathrm{euc}}}$:
@@
We retain the in‐batch contrastive retrieval losses of DrugCLIP (gao2023drugclip) and LigUnity’s listwise ranking term (feng2025foundation), applied to the hyperbolic embeddings $\tilde{\mathbf{h}}_{u}$. For each assay $A_{i}$ with query modality $u\in\{\mathrm{poc},\mathrm{seq}\}$ and its $B$ candidate ligands $\{v_{j}\}$, we compute similarity logits $s_{i,j}\;=\;\frac{1}{\tau}\,\bigl\langle\tilde{\mathbf{h}}_{u_{i}},\tilde{\mathbf{h}}_{v_{j}}\bigr\rangle.$
@@
$\mathcal{L}_{\mathrm{l}\to\mathrm{p}}^{(i)}$ $=-\frac{1}{|L_{i}|}\sum_{k\in L_{i}}\log\frac{\exp(s_{i,k})}{\sum_{n=1}^{B}\exp(s_{i,n})},$ (9)
@@
For each assay $A_{i}$ the screened ligands are sorted by measured affinity, yielding an ordered list $(v_{i,1},\dots,v_{i,B})$. Following the Plackett–Luce model 10.1145/1273496.1273513, the probability of selecting ligand $v_{i,k}$ at step $k$ (from the remaining set $\mathcal{R}_{i,k}=\{k,k+1,\dots,B\}$) is
@@ 4.4 Hyperbolic Geometry as a Structural Prior
Each pocket defines a surface of admissible directions. Its half-aperture angle is formulated by LeRPKN19; DesaiNR0V23 as
@@
Bucket 0 therefore collects the weakest binders and bucket $K$ the strongest. For every ligand we derive a bucket–specific radial limit $r_{i,j}$ and angular–scaling factor $\eta_{i,j}$
@@
where $r_{0}$ and $\eta_{0}$ are the base radius/angle for the weakest tier, and $\Delta r,\Delta\eta>0$ are the per‑tier increments. Smaller $b_{i,j}$ thus yields a smaller radius cap and a larger cone. We penalise violations in radius and angle:
@@ 4.5 Addressing Activity Cliffs with Hyperbolic Geometry
While structurally similar ligands often cluster in Euclidean space, such geometry can underrepresent functional differences—especially in activity cliffs, where minor structural changes lead to large affinity shifts. As formalized in Proposition 1, hyperbolic space provides exponentially greater separation via angular variation, offering a principled mechanism for distinguishing such cases. The theoretical derivation is provided in Appendix.
@@ 4.6 Training and Inference
$\mathcal{L}_{\text{total}}\;=$ $\underbrace{\alpha_{\mathrm{poc}}\!\Bigl(\mathcal{L}_{\text{cont}}^{\mathrm{poc}\!\leftrightarrow\!\mathrm{lig}}+\lambda_{\text{rank}}\,\mathcal{L}_{\text{rank}}^{\mathrm{poc}}\Bigr)}_{\text{pocket}\;\leftrightarrow\;\text{ligand}}$ (22)
$+\,\underbrace{\alpha_{\mathrm{seq}}\!\Bigl(\mathcal{L}_{\text{cont}}^{\mathrm{seq}\!\leftrightarrow\!\mathrm{lig}}+\lambda_{\text{rank}}\,\mathcal{L}_{\text{rank}}^{\mathrm{seq}}\Bigr)}_{\text{sequence}\;\leftrightarrow\;\text{ligand}}$
$+\,\underbrace{\gamma_{\text{cone}}\,\mathcal{L}_{\text{cone}}+\lambda_{\text{ang}}\,R_{\text{ang}}+\lambda_{\text{het}}\,R_{\text{het}}}_{\text{pocket}\;\leftrightarrow\;\text{ligand}}.$
@@ 5.1 Experimental Settings
Implemention Details. We adopt the same curated assay–level training dataset as LigUnity feng2025foundation, which is constructed from ChEMBL (mendez_chembl_2018), BindingDB 10.1093/nar/gkv1072, and PDBBind (liu2017pdbbind). For virtual screening, we strictly exclude any target UniProt IDs present in the DUD-E mysinger2012dude, LIT-PCBA tran-nguyen_lit-pcba_2020 test sets. For affinity ranking tasks, we perform ligand‐level deduplication by removing redundant small molecules and non‐redundant assay IDs. Training is run on four NVIDIA A100 GPUs for 50 epochs,, using the Adam optimizer with an initial learning rate of $1\times 10^{-4}$ and the curvature parameter $\kappa$ (absolute value of negative curvature) fixed to $1$.
Benchmark. In virtual screening, evaluations are performed on DUD-E mysinger2012dude and LIT-PCBA tran-nguyen_lit-pcba_2020. DUD-E includes 102 protein targets, each associated with experimentally verified actives and 50 property-matched decoys, designed to test enrichment capability under artificially constructed decoy scenarios. LIT-PCBA, in contrast, contains 15 targets with over 400K experimentally confirmed inactives, offering a more realistic and challenging setting without synthetic decoy bias. For affinity ranking, the evaluation is conducted on JACS wang2015accurate and Merck schindler2020large. JACS consists of eight high-quality congeneric series extracted from real lead optimization projects, emphasizing precise ranking within narrow chemical series, while Merck serves as a large-scale benchmark for FEP-based lead optimization with diverse chemical scaffolds and higher experimental noise.